The face value for WICB Limited bonds is $250,000 and has a 6 percent annual coupon. The 6 percent annual coupon bonds matures in 2035, and it is now 2020. Interest on these bonds is paid annually on December 31 of each year, and new annual coupon bonds with similar risk and maturity are currently yielding 10 percent. How much should Karen sell her bonds today?
1
Expert's answer
2020-10-29T07:06:20-0400

solution


assume we are at the beginning of 2020 and the bond matures at the end of 2035. Then


"Number\ of \ years, \ n= 16"

"Face\ value,\ F = 250,000"


"6\% \ annual\ coupon, \ C = 0.06*250,000=15,000"

"Required\ rate\ of\ return,\ r=0.10"

Present value of the bond


"PV=C*\frac{1-(1+r)^{-n}}{r} + F *(1+r)^{-n}"

"PV=15,000*\frac{1-(1.1)^{-16}}{0.1} + 250,000 *(1.1)^{-16}"

"117,355.6296+54,407.28395= 171,762.9136"

answer: the price of the bond today is $ 171,762.91